OpenStudy (bbb911):

Multiply. Write the product in its lowest term (18/7)(9/12)

OpenStudy (anonymous):

\[\frac{ 18 }{ 7 } \times \frac{ 9 }{ 12 } = \frac{ 162 }{ 84 } = \frac{ 27 }{ 14 }\]

OpenStudy (bbb911):

thank you :)